How many quarts of water does one add to 6 quarts of a 60% alcohol solution to create a 40% alcohol solution?
Answer by ptaylor(2198)

Let x=number of quarts of water to be added

The amount of pure alcohol before adding water = amount of pure alcohol after adding water. So our equation to solve is:
6(.60)=(6+x)(.40) clear parens
3.60=2.40+.4x subtract 2.40 from both sides
3.60-2.40=2.40-2.40+.40x
1.20=.40x
x=3 qts-------------------------------answer
CK
6(.6)=9(.4)
3.6=3.6

Hope this helps-----ptaylor